# Random expansions of trees with bounded height

### v1

@@ 7.5. Asymptotic elimination of aggregation functions
In this subsection we prove statement (A) above, that is, we prove that given assumptions 6.3 and 7.8 (the induction hypothesis) and the results proved earlier in Section7 we can asymptotically eliminate aggregation functions, provided that some conditions are satisfied. Some motivation for these conditions are given in the beginning of Section 8.
@@
For each $p(\bar{x},\bar{y})\in L_{0}$ define $L_{p(\bar{x},\bar{y})}$ to be the set of all $\bar{y}$-positive $\lambda(\bar{x},\bar{y})\in L_{1}$. (So $L_{p(\bar{x},\bar{y})}$ depends only on the free variables $\bar{x}$ and $\bar{y}$ but not on other properties of $p$.) Suppose that $p_{i}(\bar{x},\bar{y})\in L_{0}$ for $i=1,\ldots,k$ and let $\lambda_{i}(\bar{x},\bar{y})\in L_{p_{i}(\bar{x},\bar{y})}$ for $i=1,\ldots,k$, so $\lambda_{i}(\bar{x},\bar{y})$ is a $\bar{y}$-positive closure type over $\sigma$. Let $q_{i}(\bar{x})\in L_{0}$, $i=1,\ldots,s$, enumerate all, up to equivalence, complete closure types over $\sigma$ in the variables $\bar{x}$ and we may assume that $q_{i}(\bar{x})$ is not equivalent to $q_{j}(\bar{x})$ if $i\neq j$ (otherwise just remove some $q_{i}(\bar{x})$).
@@
Recall that we have assumed that $\mathbb{G}$ is a $PLA^{*}(\sigma)$-network of height $\rho+1$. Suppose that $\sigma\subset\sigma^{+}$ and that $\mathbb{G}^{+}$ is a $PLA^{*}(\sigma^{+})$-network of height $\rho+2$ such that $\mathbb{G}$ is the subnetwork of $\mathbb{G}^{+}$ which is induced by $\sigma$. For every $R\in\sigma^{+}\setminus\sigma$, let $\theta_{R}$ be the $PLA^{*}(\sigma)$-formula which $\mathbb{G}^{+}$ associates to $R$. Suppose that for each $R\in\sigma^{+}\setminus\sigma$, $\theta_{R}$ satisfies the assumptions on $\varphi$ in Proposition 7.26. Then Proposition 7.26 implies that part (1) of Assumption 7.8 holds if $\sigma_{\rho}$, $\sigma$ and $\mathbb{G}_{\rho}$ are replaced by $\sigma$, $\sigma^{+}$ and $\mathbb{G}$, respectively. Hence statement (A) is proved.
@@ 8. The main results
$\mathbb{P}_{n}\Big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:|\mathcal{A}(\varphi(\bar{a}))-\mathcal{A}(\bigwedge_{i=1}^{m}(\varphi_{i}(\bar{a})\to c_{i}))|\leq\varepsilon\big{\}}\Big{)}\geq 1-\varepsilon.$
@@
Recall that the aggregation functions min and max are admissible, but not continuous. Therefore Theorem 8.1 only applies to min and max if they are used together with a conditioning closure type $\chi(\bar{y},z)$ with $z$-rank 0. The next corollary states that if we strengthen the assumptions on the $PLA(\sigma)$-network $\mathbb{G}$ (thus limiting the range of distributions that such $\mathbb{G}$ can induce), then the conclusions (i) and (ii) of Theorem 8.1 hold if the formula $\varphi(\bar{x})$ uses (only) admissible aggregation functions. In particular the conclusions of the theorem hold for first-order formulas.
@@
Let $\tau$, $\sigma$, $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$, $\mathbf{W}_{n}$, $\mathbb{G}$ and $\mathbb{P}_{n}$ be as in Theorem 8.1. In particular Assumption 6.3 is adopted. Suppose that, for every $R\in\sigma\setminus\tau$, the formula $\theta_{R}$ that $\mathbb{G}$ associates to $R$ is a closure-basic formula over $\mathrm{par}(R)$ (the parents of $R$ in the underlying DAG of $\mathbb{G}$). Suppose that $\varphi(\bar{x})\in PLA^{*}(\sigma)$ and that if
@@
Proof. The conditions of Assumption 6.3 that have been removed in Assumption 8.3 are only used in the proof of Proposition 7.19 when proving the inductive step, and in this step we considered a closure type of $\bar{y}$-rank at least 2. $\square$
@@
Remark 8.8.
Let $coPLA^{*}(\sigma)$ denote the set of formulas $\varphi(\bar{x})\in PLA^{*}(\sigma)$ such that every subformula of $\varphi$ of the form (8.1) satisfies the conditions stated in Theorem 8.1. In particular, the aggregation functions maximum and minimum may only be used when the closure types $\chi_{i}(\bar{y},\bar{z})$ that we condition on have $\bar{z}$-rank 0. This intuitively means that first-order quantification can only be done over a bounded subset of a tree (a subset which is included in the closure of a tuple of vertices). So $coPLA^{*}(\sigma)$ allows only first-order quantification in a very limited way.
However in some contexts one may be more interested in the proportion (rather than existence) of elements that satisfy some (two-valued) formula. For this we can use the continuous aggregation function am which returns this proportion. If we want to express that the proportion of elements that satisfy a formula is at least as large as the proportion that satisfies another formula, then we could, tentatively, use the connective $\leq$$(x,y)$ which returns 1 if $x\leq y$ and 0 otherwise. But since this connective is not continuous it cannot be used in $PLA^{*}(\sigma)$, so instead we consider a continuous approximation of $\leq$, say $\leq_{\varepsilon}(x,y)$ = 1 if $x\leq y$, 0 if $x\geq y+\varepsilon$, and $(y+\varepsilon-x)/\varepsilon$ if $y<x<y+\varepsilon$, where $\varepsilon>0$ is small. With $\leq_{\varepsilon}$ and am we can, with formulas in $coPLA^{*}$, express approximations of statements about proportions ranging over elements that satisfy a $\bar{z}$-positive closure type over $\sigma$ for some $\bar{z}$. The idea of considering continuous approximations of connectives or aggregation functions can be used more generally to get $coPLA^{*}(\sigma)$-formulas that estimate discrete/discontinuous properties.

### v2

@@ 7.5. Asymptotic elimination of aggregation functions
In this subsection we prove statement (C) above, that is, we prove that given assumptions 6.3 and 7.8 (the induction hypothesis) and the results proved earlier in Section 7 we can asymptotically eliminate aggregation functions, provided that some conditions are satisfied. Some motivation for these conditions are given in the beginning of Section 8.
@@
For each $p(\bar{x},\bar{y})\in L_{0}$ define $L_{p(\bar{x},\bar{y})}$ to be the set of all $\bar{y}$-positive $\lambda(\bar{x},\bar{y})\in L_{1}$. (So $L_{p(\bar{x},\bar{y})}$ depends only on the free variables $\bar{x}$ and $\bar{y}$ but not on other properties of $p$.) Suppose that $p_{i}(\bar{x},\bar{y})\in L_{0}$ for $i=1,\ldots,k$ and let $\lambda_{i}(\bar{x},\bar{y})\in L_{p_{i}(\bar{x},\bar{y})}$ for $i=1,\ldots,k$, so $\lambda_{i}(\bar{x},\bar{y})$ is a $\bar{y}$-positive closure type over $\sigma$. Let $q_{i}(\bar{x})\in L_{0}$, $i=1,\ldots,s$, enumerate all, up to equivalence, complete closure types over $\sigma$ in the variables $\bar{x}$. We may assume that $q_{i}(\bar{x})$ is not equivalent to $q_{j}(\bar{x})$ if $i\neq j$.
@@
Recall that we have assumed that $\mathbb{G}$ is a $PLA^{*}(\sigma)$-network of height $\rho+1$. Suppose that $\sigma\subset\sigma^{+}$ and that $\mathbb{G}^{+}$ is a $PLA^{*}(\sigma^{+})$-network of height $\rho+2$ such that $\mathbb{G}$ is the subnetwork of $\mathbb{G}^{+}$ which is induced by $\sigma$. For every $R\in\sigma^{+}\setminus\sigma$, let $\theta_{R}$ be the $PLA^{*}(\sigma)$-formula which $\mathbb{G}^{+}$ associates to $R$. Suppose that for each $R\in\sigma^{+}\setminus\sigma$, $\theta_{R}$ satisfies the assumptions on $\varphi$ in Proposition 7.26. Then Proposition 7.26 implies that part (1) of Assumption 7.8 holds if $\sigma_{\rho}$, $\sigma$ and $\mathbb{G}_{\rho}$ are replaced by $\sigma$, $\sigma^{+}$ and $\mathbb{G}$, respectively. Hence statement (C) is proved.
@@ 8. The main results
$\mathbb{P}_{n}\Big{(}\Big{\{}\mathcal{A}\in\mathbf{W}_{n}:\Big{|}\mathcal{A}(\varphi(\bar{a}))-\mathcal{A}\Big{(}\bigwedge_{i=1}^{m}\big{(}\varphi_{i}(\bar{a})\to c_{i}\big{)}\Big{)}\Big{|}\leq\varepsilon\Big{\}}\Big{)}\geq 1-\varepsilon.$
@@
Recall that the aggregation functions min and max are admissible, but not continuous. Therefore Theorem 8.1 only applies to min and max if they are used together with a conditioning closure type $\chi(\bar{y},z)$ with $z$-rank 0. The next corollary states that if we strengthen the assumptions on the $PLA^{*}(\sigma)$-network $\mathbb{G}$ (thus limiting the range of distributions that such $\mathbb{G}$ can induce), then the conclusions (i) and (ii) of Theorem 8.1 hold if the formula $\varphi(\bar{x})$ uses (only) admissible aggregation functions. In particular the conclusions of the theorem hold for first-order formulas.
@@
Let $\tau$, $\sigma$, $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$, $\mathbf{W}_{n}$, $\mathbb{G}$ and $\mathbb{P}_{n}$ be as in Theorem 8.1. In particular Assumption 6.3 is adopted. Suppose that, for every $R\in\sigma\setminus\tau$, the formula $\theta_{R}$ that $\mathbb{G}$ associates to $R$ is a closure-basic formula over $\mathrm{par}(R)$. Suppose that $\varphi(\bar{x})\in PLA^{*}(\sigma)$ and that if
@@
Proof. The conditions of Assumption 6.3 that have been removed in Assumption 8.3 are only used in the proof of Proposition 7.19 when proving the inductive step, and the inductive step is only necessary if we only consider $p_{\tau}(\bar{x},\bar{y})$ in that proposition with $\mathrm{rank}_{\bar{y}}(p_{\tau})\geq 2$. As a consequence, all results in Section 7.4 after Proposition 7.19 restricted to closure types $p(\bar{x},\bar{y})$ with $\mathrm{rank}_{\bar{y}}(p)\leq 1$ follow from results before Proposition 7.19. Hence, Proposition 7.26 restricted to $\chi_{i}(\bar{y},\bar{z})$ with $\bar{z}$-rank at most 1 does not need Proposition 7.19. Consequently, the same holds for Corollary 8.4. $\square$
@@
Acknowledgements. The authors thank the anonymous referees for their insightful remarks and suggestions which have improved this article. Vera Koponen was partially supported by the Swedish Research Council, grant 2023-05238_VR. Vera Koponen and Yasmin Tousinejad were partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.